OpenStudy (anonymous):

Simplify the following expression: 3(3z - 4 ) - (20 z - 5)

OpenStudy (anonymous):

what is the equivalent? if it is 3(3z - 4 ) - (20 z - 5)= 0 then 9z -12 -20z +5 =0 -11z = 7 z= -7/11

OpenStudy (anonymous):

I'll help ya buddy!

OpenStudy (anonymous):

Is this what your looking for 3(3z - 4 ) - (20 z - 5)